# Relation between generalized and ordinary cluster algebras

### v1

@@ 1. Introduction
Cluster algebras were introduced by Fomin and Zelevinsky [FZ02]. They naturally appear in several areas of mathematics and have been studied and developed by many authors. In [CS14], Chekhov and Shapiro introduced a natural generalization of cluster algebras called generalized cluster algebras (GCAs) by replacing the binomial exchange polynomials of (ordinary) cluster algebras (CAs) with the multinomial ones. It turned out that GCAs share several important properties with CAs, for example, the Laurent phenomenon [CS14], the separation formulas [Nak15a], quantizations [Nak15b, BCMX18, FPY24], scattering diagram formulation [CKM21, Mou24], the Laurent positivity [Mou24, BLM25a, BLM25b], etc.
@@
Due to the importance of the above result, in this paper we apply the same idea and method to GCAs with coefficients in an arbitrary semifield and obtain a parallel result for them. To do it, in contrast to [RW25], we concentrate on the class of GCAs with $y$-variables in [Nak15a, NR16, Nak24], which is parallel to the one of CAs in [FZ07]. Compared with the formulation of GCAs in [RW25], the advantage of ours is two-folded. Firstly, we are away from the specific problem arising from the exchange matrix presentation of tropical coefficients in [RW25]. (A considerable part of the paper [RW25] is devoted to resolve the issues related to it.) Secondly, the parallelism of the factorization mechanism between cluster variables ($x$-variables) and coefficients ($y$-variables) is transparent. We also present the relations between the $C$-matrices, the $G$-matrices, and the $F$-polynomials of a generalized cluster pattern and those of the corresponding composite cluster pattern. As the most important corollary of our result (Theorem 5.4), we obtain the Laurent positivity of the above GCAs from the one of CAs [GHKK18]. This property was proved by [BLM25b] with the scattering diagram method only recently.
@@
We thank Zhichao Chen and Yasuaki Gyoda for discussions. This work is supported in part by JSPS grant No. JP22H01114 and No. JP25KJ1438. Ryota Akagi is also supported by Chubei Itoh Foundation.
@@ 2. Generalized cluster algebras with $y$-variables
In this section, we recall the definition of generalized cluster algebras (GCAs) with $y$-variables following [Nak15a, NR16, Nak24].
@@ 5.1. Realization of generalized cluster patterns
As an immediate consequence of Theorem 5.4, the Laurent positivity of ${\bf\Sigma}^{\mathrm{g}}$ follows from the one of the corresponding ordinary cluster pattern $\bf\Sigma$}.
@@ 6.1. Separation formulas
In parallel to the ordinary case [FZ07], to any generalized cluster pattern $\mathbf{\Sigma}^{\mathrm{g}}$ with $y$-variables, one can associate the collections of the $C$-matrices $\{C^{\mathrm{g}}_{t}\}_{t\in\mathbb{T}_{n}}$, the $G$-matrices $\{G^{\mathrm{g}}_{t}\}_{t\in\mathbb{T}_{n}}$, and the $F$-polynomials $\{F^{\mathrm{g}}_{i;t}\}_{i=1,\dots,n;t\in\mathbb{T}_{n}}$ [Nak15a]. Here, each $F$-polynomial $F^{\mathrm{g}}_{i;t}(\mathbf{y},\mathbf{z})$ is a polynomial in formal variables $\mathbf{y}=(y_{i})_{i=1}^{n}$ and $\mathbf{z}=(z_{il})_{(i,l)\in\mathcal{D}^{\prime}}$ with (nonnegative) integer coefficients, where $\mathcal{D}^{\prime}=\{(i,l)\mid i=1,\dots,n;\ l=1,\dots,r_{i}-1\}$. Then, the $x$- and $y$-variables of $\mathbf{\Sigma}^{\mathrm{g}}$ are expressed by the following separation formulas [Nak15a, NR16]:
@@ 6.2. Relations for $C$-matrices and $G$-matrices
For any $i,j=1,\dots,n$, set $\delta_{ij}=1$ if $i=j$ and $\delta_{ij}=0$ if $i\neq j$. The $C$-matrices and the $G$-matrices are determined by the initial condition $c_{ij;t_{0}}^{\mathrm{g}}=g_{ij;t_{0}}^{\mathrm{g}}=\delta_{ij}$ and the following recursions for each pair of $k$-adjacent vertices $t,t^{\prime}\in\mathbb{T}_{n}$ [Nak15a, NR16]:
@@ 6.3. Relation for $F$-polynomials
and we recursively define $\mathcal{Z}_{i;t}(u)$ ($t\in\mathbb{T}_{n}$) by $\mathcal{Z}_{i;t^{\prime}}(u)=\mathcal{Z}_{i;t}(u)$ if $i\neq k$ and $\mathcal{Z}_{k;t^{\prime}}(u)=\mathcal{Z}_{k;t}^{\mathrm{rec}}(u)$ for any pair of $k$-adjacent vertices $t,t^{\prime}\in\mathbb{T}_{n}$. The $F$-polynomials are determined by the initial condition $F_{i;t_{0}}=1$ ($i=1,\dots,n$) and the following recursion for each pair of $k$-adjacent vertices $t,t^{\prime}\in\mathbb{T}_{n}$ [Nak15a, NR16]:

### v2

@@ 1. Introduction
Cluster algebras were introduced by Fomin and Zelevinsky [FZ02]. They naturally appear in several areas of mathematics and have been studied and developed by many authors. In [CS14], Chekhov and Shapiro introduced a natural generalization of cluster algebras called generalized cluster algebras (GCAs) by replacing the binomial exchange polynomials of (ordinary) cluster algebras (CAs) with the multinomial ones. It turned out that GCAs share several important properties with CAs, for example, the Laurent phenomenon [CS14], the separation formulas [NAK15b], quantizations [NAK15a, BCD+18, FPY24], scattering diagram formulation [CKM21, MOU24], the Laurent positivity [MOU24, BLM25b, BLM25a], etc.
@@
Due to the importance of the above result, in this paper we apply the same idea and method to GCAs with coefficients in an arbitrary semifield and obtain a parallel result for them. To do it, in contrast to [RW25], we concentrate on the class of GCAs with $y$-variables in [NAK15b, NR16, NAK24], which is parallel to the one of CAs in [FZ07]. Compared with the formulation of GCAs in [RW25], the advantage of ours is two-folded. Firstly, we are away from the specific problem arising from the exchange matrix presentation of tropical coefficients in [RW25]. (A considerable part of the paper [RW25] is devoted to resolve the issues related to it.) Secondly, the parallelism of the factorization mechanism between cluster variables ($x$-variables) and coefficients ($y$-variables) is transparent. We also present the relations between the $C$-matrices, the $G$-matrices, and the $F$-polynomials of a generalized cluster pattern and those of the corresponding composite cluster pattern.
@@
We thank Zhichao Chen and Yasuaki Gyoda for discussions. We thank Woojin Choi for pointing out the error in the early version of the manuscript. This work is supported in part by JSPS grant No. JP22H01114 and No. JP25KJ1438. Ryota Akagi is also supported by Chubei Itoh Foundation.
@@ 2. Generalized cluster algebras with $y$-variables
In this section, we recall the definition of generalized cluster algebras (GCAs) with $y$-variables following [NAK15b, NR16, NAK24].
@@ 6.1. Separation formulas
In parallel to the ordinary case [FZ07], to any generalized cluster pattern $\mathbf{\Sigma}^{\mathrm{g}}$ with $y$-variables, one can associate the collections of the $C$-matrices $\{C^{\mathrm{g}}_{t}\}_{t\in\mathbb{T}_{n}}$, the $G$-matrices $\{G^{\mathrm{g}}_{t}\}_{t\in\mathbb{T}_{n}}$, and the $F$-polynomials $\{F^{\mathrm{g}}_{i;t}\}_{i=1,\dots,n;t\in\mathbb{T}_{n}}$ [NAK15b]. Here, each $F$-polynomial $F^{\mathrm{g}}_{i;t}(\mathbf{y},\mathbf{z})$ is a polynomial in formal variables $\mathbf{y}=(y_{i})_{i=1}^{n}$ and $\mathbf{z}=(z_{il})_{(i,l)\in\mathcal{D}^{\prime}}$ with (nonnegative) integer coefficients, where $\mathcal{D}^{\prime}=\{(i,l)\mid i=1,\dots,n;\ l=1,\dots,r_{i}-1\}$. Then, the $x$- and $y$-variables of $\mathbf{\Sigma}^{\mathrm{g}}$ are expressed by the following separation formulas [NAK15b, NR16]:
@@ 6.2. Relations for $C$-matrices and $G$-matrices
For any $i,j=1,\dots,n$, set $\delta_{ij}=1$ if $i=j$ and $\delta_{ij}=0$ if $i\neq j$. The $C$-matrices and the $G$-matrices are determined by the initial condition $c_{ij;t_{0}}^{\mathrm{g}}=g_{ij;t_{0}}^{\mathrm{g}}=\delta_{ij}$ and the following recursions for each pair of $k$-adjacent vertices $t,t^{\prime}\in\mathbb{T}_{n}$ [NAK15b, NR16]:
@@ 6.3. Relation for $F$-polynomials
and we recursively define $\mathcal{Z}_{i;t}(u)$ ($t\in\mathbb{T}_{n}$) by $\mathcal{Z}_{i;t^{\prime}}(u)=\mathcal{Z}_{i;t}(u)$ if $i\neq k$ and $\mathcal{Z}_{k;t^{\prime}}(u)=\mathcal{Z}_{k;t}^{\mathrm{rec}}(u)$ for any pair of $k$-adjacent vertices $t,t^{\prime}\in\mathbb{T}_{n}$. The $F$-polynomials are determined by the initial condition $F_{i;t_{0}}=1$ ($i=1,\dots,n$) and the following recursion for each pair of $k$-adjacent vertices $t,t^{\prime}\in\mathbb{T}_{n}$ [NAK15b, NR16]:
@@
Remark 6.5.
The equality (6.23) does not directly imply the positivity of the coefficients in $F_{i;t}^{\mathrm{g}}$ (the Laurent positivity of $x_{i;t}^{\mathrm{g}}$ proved by [BLM25a]) from the one of $F_{il;t}^{\mathrm{c}}$ due to possible signs such as $s_{1}^{2}+s_{2}^{2}=(s_{1}+s_{2})^{2}-2s_{1}s_{2}$.